-7+3,
-24+(-17),
11+
(-11),
-3+14 ,
-3.6+1.9,
-15+(-7),
-18+18 ,
-3.8+(-9.4),
-3/5+2/5,
-4/9+(-6/9),
35+(-14)+(-19)+(-5),
28+(-44)+17+31+(-94),
-44+(3/8)+95+(5/8),
24+3.1+(-44)+(8.2)+63.
1
Expert's answer
2011-07-27T05:08:03-0400
-7+3 = -4,
-24+(-17) = -24-17 = -41,
11+
(-11) = 11 -11 = 0,
-3+14 = 14-3 = 11,
-3.6+1.9 = -1.7,
-15+(-7) = -15 - 7 = -22,
-18+18 = 0,
-3.8+(-9.4) = - 3.8 - 9.4 = - 13.2,
-3/5+2/5 = (-3 +2)/5 = -1/5,
-4/9+(-6/9) = (-4 -6)/9 = -10/9,
35+(-14)+(-19)+(-5) = 35 - 14 - 19 - 5 = -3,
28+(-44)+17+31+(-94) = 28 - 44 + 17 +31 - 94 = - 62,
-44+(3/8)+95+(5/8) = -44 + 95 + (3+5)/8 = 51 + 8/8 = 51+ 1 = 52,
24+3.1+(-44)+(8.2)+63 = 24 + 3.1 - 44 + 8.2 + 63 = 54.3.